OpenStudy (anonymous):

what are the period and phase shift for f(x)=3 tan(4x+pi)?

OpenStudy (anonymous):

period of tangent is \(\pi\) period of \(\tan(bx)\) is \(\frac{\pi}{b}\) in your case \(b=4\)